# Forecast constraints on the axion-photon coupling from interstellar medium heating

### v1

@@ Abstract
In interstellar media characterized by a nonrelativistic plasma of electrons and heavy ions, we study the effect of axion dark matter coupled to photons on the dynamics of an electric field. In particular, we assume the presence of a background magnetic field aligned in a specific direction. We show that there is an energy transfer from the oscillating axion field to photons and then to the plasma induced by forced resonance. This resonance is most prominent for the axion mass $m_{\phi}$ equivalent to the plasma frequency $\omega_{p}$. Requiring that the heating rate of the interstellar medium caused by the energy transfer does not exceed the observed astrophysical cooling rate, we place forecast constraints on the axion-photon coupling $g$ for several different amplitudes of the background magnetic field $B_{0}$. By choosing a typical value $B_{0}=10^{-5}$ G, we find that, for the resonance mass $m_{\phi}=\omega_{p}$, the upper limit of $g$ can be stronger than those derived from other measurements in the literature. With increased values of $B_{0}$, it is possible to put more stringent constraints on $g$ for a wider range of the axion mass away from the resonance point.
@@ I Introduction
An axion $\phi$ may interact with photons through the Chern-Simons coupling ${\cal L}=-g\phi F_{\mu\nu}\widetilde{F}^{\mu\nu}/4$, where $g$ is a coupling constant and $F_{\mu\nu}$ is an electromagnetic field strength with $\widetilde{F}^{\mu\nu}$ its Hodge dual. In terms of the electric field ${\bm{E}}$ and the magnetic field ${\bm{B}}$, this interaction can be interpreted as the inner product $-g\phi{\bm{E}}\cdot{\bm{B}}$. Since the conversion of axions to photons can occur in the presence of external magnetic fields [26], the laboratory experiments such as a “light-shining-through-walls” measurement [27] put upper limits on the coupling $g$. For the mass range $m_{\phi}\lesssim 10^{-4}$ eV, the ALPS collaboration has constrained $g\leq 5\times 10^{-8}$ GeV-1 [28]. By considering axions that are produced in the Sun and convert to $X$-rays in the magnetic field at the detector, the CAST [29] and IAXO [30] experiments placed the limit $g\leq 4\times 10^{-12}$ GeV-1 for $m_{\phi}\leq 5\times 10^{-3}$ eV.
@@
In this paper, we will propose a new method of probing the axion-photon coupling through the observed cooling rate of interstellar media. We consider a nonrelativistic plasma of electrons and heavy ions with the electron number density $n_{e}$. In the absence of the axion-photon coupling, the collective movement of electrons gives rise to the plasma oscillation of an electric field ${\bm{E}}$ with the angular frequency $\omega_{p}=\sqrt{n_{e}e^{2}/m_{e}}\simeq 10^{-11}\,{\rm eV}\,(n_{e}/0.1~{}{\rm cm}^{-3})^{1/2}$, where $e$ is the elementary charge and $m_{e}$ is the electron mass [47, 48]. For the typical number density $n_{e}\approx 0.1$ cm-3 of dilute plasmas in interstellar media, the plasma frequency $\omega_{p}$ is of order $10^{-11}$ eV. The existence of heavy ions works to generate friction for the dynamics of electrons characterized by the damping rate $\nu$. Since we typically have $\nu\ll\omega_{p}$ for the interstellar plasma, the electric field exhibits damped oscillations with the decay time scale $\nu^{-1}$, which is much longer than the oscillating time scale $\omega_{p}^{-1}$.
@@
If there are background magnetic fields in the interstellar region, the interaction ${\cal L}=-g\phi F_{\mu\nu}\widetilde{F}^{\mu\nu}/4$ can lead to the energy conversion from axions to photons in the plasma. The observations of interstellar magnetic fields in the galactic center suggest that their amplitudes may span in the range $B_{0}=10^{-5}~{}{\rm G}\sim 10^{-3}$ G [49]. In this paper, we will consider an external magnetic field ${\bm{B}}_{0}$ aligned in a specific direction and compute the heating rate of interstellar media induced by the axion-photon coupling. For this purpose, we deal with the axion as nonrelativistic dark matter oscillating around a minimum of the potential $V(\phi)=m_{\phi}^{2}\phi^{2}/2$. In the two-dimensional plane perpendicular to the direction of ${\bm{B}}_{0}$, the electric-field components still exhibit damped elliptic motions combined with the plasma oscillation and the cyclotron orbit in the normal manner. Along the direction of ${\bm{B}}_{0}$, however, we will show that the energy density of axions is converted to that of photons through forced resonance. In particular, the energy transfer rate $\dot{Q}$ is largest for the resonance mass $m_{\phi}$ equivalent to the plasma frequency $\omega_{p}$.
@@
The observations of interstellar media like Leo T have constrained the cooling rate $\dot{C}$ as well as $\omega_{p}$ and $\nu$ [50, 51, 52, 53]. These astrophysical data can be used to probe the properties of dark matter [54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. In our case, the heating rate $\dot{Q}$ induced by the axion-photon coupling should not exceed the observed cooling rate, which translates to the upper limit of $g$. At the resonance point, we will derive an analytic formula for the bound on $g$. For the plasma frequency in the range $10^{-12}~{}{\rm eV}\lesssim\omega_{p}\lesssim 10^{-10}~{}{\rm eV}$, the upper limit of $g$ at the resonance mass $m_{\phi}=\omega_{p}$ can be smaller than the MWD bound mentioned above by choosing the typical galactic magnetic field strength $B_{0}\approx 10^{-5}$ eV. In particular, the constraint on $g$ at $m_{\phi}=\omega_{p}$ tends to be stronger for smaller $\nu$ due to the occurrence of sharper resonance. With the increase of $B_{0}$, we can place tighter bounds on $g$ for a broader mass range away from the resonance point. While the precise magnetic field strengths have been unknown for interstellar media with observed cooling rates, upcoming observations may be able to put interesting constraints on $g$ for the mass range $10^{-15}~{}{\rm eV}\lesssim m_{\phi}\lesssim 10^{-7}~{}{\rm eV}$.
@@ II Plasma in interstellar media
$\nu=\frac{4\sqrt{2\pi}\alpha^{2}n_{e}}{6(m_{e}T^{3})^{1/2}}\ln\Lambda_{c}\simeq 1.2\times 10^{-22}~{}{\rm eV}\left(\frac{n_{e}}{0.1~{}{\rm cm}^{-3}}\right)\left(\frac{10^{4}~{}{\rm K}}{T}\right)^{3/2}\ln\Lambda_{c}\,,$ (3)
@@
$\Lambda_{c}=\frac{4\pi T^{3}}{\alpha^{3}n_{e}}\simeq 2.7\times 10^{22}\left(\frac{0.1~{}{\rm cm}^{-3}}{n_{e}}\right)\left(\frac{T}{10^{4}~{}{\rm K}}\right)^{3}\,.$ (4)
@@
$\omega_{p}=\sqrt{\frac{n_{e}e^{2}}{m_{e}}}\simeq 1.2\times 10^{-11}~{}{\rm eV}\left(\frac{n_{e}}{0.1~{}{\rm cm}^{-3}}\right)^{1/2}$ (7)
is the plasma frequency.
In Table 1, we summarise the observed gas-rich astrophysical media and their associated values of $n_{e}$ and $T$. The plasma frequency $\omega_{p}$ and the electron-ion friction constant $\nu$ are computed according to the relations (7) and (3), respectively. It follows that $\omega_{p}\gg\nu$ for all these interstellar media. In the last column of Table 1, we also show the astrophysical cooling rate of the gas $\dot{C}$ for each medium. It is possible to exploit these observed values of $\dot{C}$ to put constraints on the coupling $g$ between axions and electromagnetic fields.
Interstellar media $n_{e}$ (cm-3) $T$ (K) $\omega_{p}$ (eV) $\nu$ (eV) $\dot{C}$ (erg cm-3 s-1)
WNM (Leo T) 0.06 6100 $9.1\times 10^{-12}$ $7.6\times 10^{-21}$ $7\times 10^{-30}$
CNM (G33.4--8.0) 0.4 400 $2.3\times 10^{-11}$ $2.4\times 10^{-18}$ $1.46\times 10^{-27}$
MC (MW) 100 50 $3.7\times 10^{-10}$ $9.8\times 10^{-15}$ $3.16\times 10^{-24}$
CNM (MW) 30 100 $2.0\times 10^{-10}$ $1.2\times 10^{-15}$ $2.02\times 10^{-24}$
WNM (MW) 0.6 5000 $2.9\times 10^{-11}$ $9.7\times 10^{-20}$ $8.3\times 10^{-27}$
WIM (MW) 0.3 $10^{4}$ $2.0\times 10^{-11}$ $1.8\times 10^{-20}$ $5.4\times 10^{-26}$
HIM (MW) 0.003 106 $2.0\times 10^{-12}$ $2.5\times 10^{-25}$ $1.3\times 10^{-27}$
Table 1: Interstellar media, their electron number densities $n_{e}$, temperatures $T$, plasma frequencies $\omega_{p}$, electron-ion friction constants $\nu$, and cooling rates $\dot{C}$. The different phases are labeled as MC (molecular clouds), CNM (cold neutral medium), WNM (warm neutral medium),WIM (warm ionized medium), and HIM (hot ionized medium). MW denotes the data in the Milky Way and Leo T is a gas-rich dwarf galaxy. We take the observed values of $n_{e}$, $T$, and $\dot{C}$ from Ref. [53].
@@ III Axions coupled to electromagnetic fields
As a warm-up, let us revisit the case where ${\bm{B}}={\bm{0}}$. Under the condition (23), we ignore the term $k^{2}{\bm{A}}$ relative to $\ddot{\bm{A}}$ in Eq. (22). On account of Eq. (24), we have ${\bm{J}}\simeq-\dot{\bm{E}}$ and $\dot{\bm{J}}\simeq\omega_{p}^{2}{\bm{E}}-\nu{\bm{J}}$ from Eqs. (22) and (25), respectively. Combing these two equations leads to
@@
$\omega_{c}=1.2\times 10^{-13}~{}{\rm eV}\left(\frac{B_{0}}{10^{-5}~{}{\rm G}}\right)\,,$ (36)
where we used the conversion of unit $1\,{\rm G}=1.95\times 10^{-2}$ eV2. For a typical magnetic field strength $B_{0}\approx 10^{-5}$ G, we have $\omega_{c}\simeq 10^{-13}$ eV. In such a case, so long as $n_{e}$ is not much less than the order $0.1$ cm-3, the plasma frequency (7) is much larger than $\omega_{c}$, i.e., $\omega_{p}\gg\omega_{c}$. For $B_{0}\gtrsim 10^{-3}$ G, we have the opposite inequality $\omega_{p}\lesssim\omega_{c}$.
@@
The amplitudes of $A_{x}$ and $E_{x}$ blow up around the axion mass $m_{\phi}=\omega_{p}$, which is a characteristic feature of forced resonance. At $m_{\phi}=\omega_{p}$, the squared amplitude of $A_{x}$ has a peak value ${\cal D}_{\rm M}^{2}=(g\bar{\phi}_{0}B_{0})^{2}(\omega_{p}^{2}+\nu^{2})/(\omega_{p}^{2}\nu^{2})$. The region in which ${\cal D}^{2}\geq{\cal D}_{\rm M}^{2}/2$ corresponds to $\omega_{p}-\nu/2\leq m_{\phi}\leq\omega_{p}+\nu/2$, so that the resonance width is given by $\Delta m_{\phi}=\nu$. Since we are now considering the case $\nu\ll\omega_{p}$, ${\cal D}^{2}$ has a sharp peak around $m_{\phi}=\omega_{p}\pm\nu/2$.
@@ IV Heating rate induced by forced resonance
Taking the homogenous limit $k\to 0$ in Eq. (37) and using the property $E_{x}\simeq-\dot{A}_{x}$, we have
@@
We note that this has been derived by taking the large-scale limit $k\to 0$ in Fourier space. This amounts to neglecting all the spatial derivatives in real-space Eqs. (12)-(14). Then, we can interpret $\rho_{x}=E_{x}^{2}/2+\omega_{p}^{2}A_{x}^{2}/2$ as the $x$-component of the photon energy density in real space obtained by keeping the time derivatives alone. The temporal change of $\rho_{x}$ averaged over the oscillating period of the axion $T_{\phi}=2\pi/m_{\phi}$ is given by
@@
$g\leq 1.9\times 10^{-14}~{}{\rm GeV}^{-1}\left(\frac{B_{0}}{10^{-5}~{}{\rm G}}\right)^{-1}\left(\frac{\nu}{10^{-20}~{}{\rm eV}}\right)^{1/2}\left(\frac{\rho_{\rm DM}}{0.3~{}{\rm GeV}\,{\rm cm}^{-3}}\right)^{-1/2}\left(\frac{\dot{C}}{10^{-27}~{}{\rm erg}~{}{\rm cm}^{-3}~{}{\rm s}^{-1}}\right)^{1/2}\,,$ (58)
which is valid at the resonance point with the axion mass $m_{\phi}=\omega_{p}$. For $B_{0}=10^{-5}$ G, $\nu=10^{-20}$ eV, $\rho_{\rm DM}=0.3$ GeV cm-3, and $\dot{C}=10^{-27}~{}{\rm erg}~{}{\rm cm}^{-3}~{}{\rm s}^{-1}$, we have the forecast constraint $g\leq 1.9\times 10^{-14}$ GeV-1. This is tighter than the observational bound $g\leq 5.4\times 10^{-12}$ GeV-1 extracted from polarization measurements of the MWD stars for the mass range $m_{\phi}\leq 3\times 10^{-7}$ eV [35]. Should we obtain tighter observational bounds on the cooling rate $\dot{C}$, the constraint on $g$ becomes more stringent. For larger $B_{0}$, the upper limit on $g$ also decreases in proportion to $B_{0}^{-1}$.
Figure 1: The coupling constant $g$ versus the axion mass $m_{\phi}$ for $\dot{Q}$ equivalent to the two cooling rates $\dot{C}=10^{-27}$ erg cm-3 s-1 (left) and $\dot{C}=10^{-31}$ erg cm-3 s-1 (right). In the left and right panels, we choose the background magnetic field strength to be $B_{0}=10^{-5}$ G and $B_{0}=1$ G, respectively, with $\omega_{p}=10^{-11}$ eV and $\rho_{\rm DM}=0.3\,{\rm GeV}$ cm-3 in both cases. The thick colored lines correspond to (a) $\nu=10^{-13}\omega_{p}$, (b) $\nu=10^{-9}\omega_{p}$, and (c) $\nu=10^{-5}\omega_{p}$, respectively. For decreasing $\nu$, the sharper resonance occurs with smaller minimum values of $g$. The dashed black lines represent the observational bound $g\leq 5.4\times 10^{-12}$ GeV-1 derived from polarization measurements of the MWD stars for the axion mass range $m_{\phi}\leq 3\times 10^{-7}$ eV [35].
@@
In Fig. 1, we plot $g_{\rm max}$ versus $m_{\phi}$ for three different values of $\nu$. We choose two combinations of $\dot{C}$ and $B_{0}$ in the left and right panels. In both panels, the plasma frequency and the dark matter density are fixed to be $\omega_{p}=10^{-11}$ eV and $\rho_{\rm DM}=0.3\,{\rm GeV}$ cm-3. For a given $\nu$, $g_{\rm max}$ has a minimum at the resonance mass $m_{\phi}=\omega_{p}$. For decreasing $\nu$, the minimum values of $g_{\rm max}$ get smaller as expected. In both the two asymptotic mass regions $m_{\phi}\ll\omega_{p}$ and $m_{\phi}\gg\omega_{p}$, we can confirm that $g_{\rm max}$ tends to increase for smaller $\nu$.

### v2

@@ Abstract
In interstellar media characterized by a nonrelativistic plasma of electrons and heavy ions, we study the effect of axion dark matter coupled to photons on the dynamics of an electric field. In particular, we assume the presence of a background magnetic field aligned in a specific direction. We show that there is an energy transfer from the oscillating axion field to photons and then to the plasma induced by forced resonance. This resonance is most prominent for the axion mass $m_{\phi}$ equivalent to the plasma frequency $\omega_{p}$. Requiring that the heating rate of the interstellar medium caused by the energy transfer does not exceed the observed astrophysical cooling rate, we place forecast constraints on the axion-photon coupling $g$ for several different amplitudes of the background magnetic field $B_{0}$. By choosing a typical value $B_{0}=10^{-6}$ G, we find that, for the resonance mass $m_{\phi}=\omega_{p}$, the upper limit of $g$ can be stronger than those derived from other measurements in the literature. With increased values of $B_{0}$, it is possible to put more stringent constraints on $g$ for a wider range of the axion mass away from the resonance point.
@@ I Introduction
An axion $\phi$ may interact with photons through the Chern-Simons coupling ${\cal L}=-g\phi F_{\mu\nu}\widetilde{F}^{\mu\nu}/4$, where $g$ is a coupling constant and $F_{\mu\nu}$ is an electromagnetic field strength with $\widetilde{F}^{\mu\nu}$ its Hodge dual. In terms of the electric field ${\bm{E}}$ and the magnetic field ${\bm{B}}$, this interaction can be interpreted as the inner product $-g\phi{\bm{E}}\cdot{\bm{B}}$. Since the conversion of axions to photons can occur in the presence of external magnetic fields [26], the laboratory experiments such as a “light-shining-through-walls” measurement [27] put upper limits on the coupling $g$. For the mass range $m_{\phi}\lesssim 10^{-4}$ eV, the ALPS collaboration has constrained $g\leq 5\times 10^{-8}$ GeV-1 [28]. By considering axions produced in the Sun that convert into X-rays within a laboratory magnetic field, the CAST experiment [29] has already placed a limit $g\leq 0.66\times 10^{-10}~{}\text{GeV}^{-1}$ for $m_{\phi}\leq 2\times 10^{-2}~{}\text{eV}$. In contrast, the proposed IAXO experiment [30] has not obtained the data yet but is designed to increase the signal-to-noise ratio by approximately 4-5 orders of magnitude compared to CAST.
@@
In this paper, we will propose a new method of probing the axion-photon coupling through the observed cooling rate of interstellar media. We consider a nonrelativistic plasma of electrons and heavy ions with the electron number density $n_{e}$. In the absence of the axion-photon coupling, the collective movement of electrons gives rise to the plasma oscillation of an electric field ${\bm{E}}$ with the angular frequency $\omega_{p}=\sqrt{n_{e}e^{2}/m_{e}}\simeq 10^{-12}\,{\rm eV}\,(n_{e}/10^{-3}~{}{\rm cm}^{-3})^{1/2}$, where $e$ is the elementary charge and $m_{e}$ is the electron mass [47, 48]. For the typical number density $n_{e}\approx 10^{-3}$ cm-3 of dilute plasmas in the interstellar media, the plasma frequency $\omega_{p}$ is of order $10^{-12}$ eV. The existence of heavy ions works to generate friction for the dynamics of electrons characterized by the damping rate $\nu$. Since we typically have $\nu\ll\omega_{p}$ for the interstellar plasma, the electric field exhibits damped oscillations with the decay time scale $\nu^{-1}$, which is much longer than the oscillating time scale $\omega_{p}^{-1}$. If there are background magnetic fields in the interstellar region, the interaction ${\cal L}=-g\phi F_{\mu\nu}\widetilde{F}^{\mu\nu}/4$ can lead to the energy conversion from axions to photons in the plasma.
@@
A typical value of the galactic magnetic field in the Milky Way has been estimated to be of order $B_{0}=10^{-6}$ G [49, 50]. The observations of interstellar magnetic fields in the galactic center suggest that their amplitudes may span in the range $B_{0}=10^{-5}~{}{\rm G}\sim 10^{-3}$ G [51]. In this paper, we will consider an external magnetic field ${\bm{B}}_{0}$ aligned in a specific direction and compute the heating rate of interstellar media induced by the axion-photon coupling. For this purpose, we deal with the axion as nonrelativistic dark matter oscillating around a minimum of the potential $V(\phi)=m_{\phi}^{2}\phi^{2}/2$. In the two-dimensional plane perpendicular to the direction of ${\bm{B}}_{0}$, the electric-field components still exhibit damped elliptic motions combined with the plasma oscillation and the cyclotron orbit in the normal manner. Along the direction of ${\bm{B}}_{0}$, however, we will show that the energy density of axions is converted to that of photons through forced resonance. In particular, the energy transfer rate $\dot{Q}$ is largest for the resonance mass $m_{\phi}$ equivalent to the plasma frequency $\omega_{p}$.
@@
The observations of interstellar media like Leo T have constrained the cooling rate $\dot{C}$ and the temperature $T$ [52, 53, 54, 55]. They can be used to probe the properties of dark matter [56, 57, 58, 59, 60, 61, 62, 50, 63, 64]. In our case, the heating rate $\dot{Q}$ induced by the axion-photon coupling should not exceed the observed cooling rate, which translates to the upper limit of $g$. At the resonance point, we will derive an analytic formula for the bound on $g$. For the plasma frequency of order $10^{-12}~{}{\rm eV}$, the upper limit of $g$ at the resonance mass $m_{\phi}=\omega_{p}$ can be smaller than the MWD bound mentioned above by choosing the typical galactic magnetic field strength $B_{0}\approx 10^{-6}$ eV. In particular, the constraint on $g$ at $m_{\phi}=\omega_{p}$ tends to be stronger for smaller $\nu$ due to the occurrence of sharper resonance. With the increase of $B_{0}$, we can place tighter bounds on $g$ for a broader mass range away from the resonance point. While the precise magnetic field strengths have been unknown for interstellar media with observed cooling rates, upcoming observations may be able to put interesting constraints on $g$ for the mass range $10^{-15}~{}{\rm eV}\lesssim m_{\phi}\lesssim 10^{-9}~{}{\rm eV}$.
@@ II Plasma in interstellar media
$\nu=\frac{4\sqrt{2\pi}\alpha^{2}n_{e}}{6(m_{e}T^{3})^{1/2}}\ln\Lambda_{c}\simeq 1.2\times 10^{-24}~{}{\rm eV}\left(\frac{n_{e}}{10^{-3}~{}{\rm cm}^{-3}}\right)\left(\frac{10^{4}~{}{\rm K}}{T}\right)^{3/2}\ln\Lambda_{c}\,,$ (3)
@@
$\Lambda_{c}=\frac{4\pi T^{3}}{\alpha^{3}n_{e}}\simeq 2.7\times 10^{24}\left(\frac{10^{-3}~{}{\rm cm}^{-3}}{n_{e}}\right)\left(\frac{T}{10^{4}~{}{\rm K}}\right)^{3}\,.$ (4)
@@
$\omega_{p}=\sqrt{\frac{n_{e}e^{2}}{m_{e}}}\simeq 1.2\times 10^{-12}~{}{\rm eV}\left(\frac{n_{e}}{10^{-3}~{}{\rm cm}^{-3}}\right)^{1/2}$ (7)
is the plasma frequency. A dwarf galaxy Leo T has the electron number density $n_{e}\approx 10^{-3}$ cm-3 in the warm neutral region [55], in which case $\omega_{p}$ is of order $10^{-12}$ eV. In the Milky Way, there is a model for the free electron density called NE2001, which indicates that $n_{e}$ spans in the range $10^{-4}~{}{\rm cm}^{-3}\lesssim n_{e}\lesssim 0.1~{}{\rm cm}^{-3}$ depending on the region inside the galaxy [65]. In the vicinity of the Solar system, the electron number density is $n_{e}\approx 0.04$ cm-3, whose value is chosen to estimate $\omega_{p}$ in Ref. [57]. In this paper, we choose $n_{e}=10^{-3}$ cm-3 as the typical electron number density.
In Table 1, we summarise the observed gas-rich astrophysical media, their temperatures, and the astrophysical cooling rate of the gas $\dot{C}$ for each medium. The electron-ion friction constant $\nu$ can be computed according to the relation (3) with (4). For $n_{e}=10^{-3}$ cm-3, we have $\nu\simeq 1.4\times 10^{-22}$ eV for WNM (Leo T), which is much smaller than $\omega_{p}\simeq 1.2\times 10^{-12}$ eV. The property $\nu\ll\omega_{p}$ also holds for the other interstellar media shown in Table 1. By using the observed values of $\dot{C}$, it is possible to put constraints on the coupling $g$ between axions and electromagnetic fields, which will be addressed in the subsequent sections.
Interstellar media $T$ (K) $\dot{C}$ (erg cm-3 s-1)
WNM (Leo T) 6100 $7\times 10^{-30}$
CNM (G33.4--8.0) 400 $1.46\times 10^{-27}$
MC (MW) 50 $3.16\times 10^{-24}$
CNM (MW) 100 $2.02\times 10^{-24}$
WNM (MW) 5000 $8.3\times 10^{-27}$
WIM (MW) $10^{4}$ $5.4\times 10^{-26}$
HIM (MW) $10^{6}$ $1.3\times 10^{-27}$
Table 1: Interstellar media, their temperatures $T$ and cooling rates $\dot{C}$. The different phases are labeled as MC (molecular clouds), CNM (cold neutral medium), WNM (warm neutral medium), WIM (warm ionized medium), and HIM (hot ionized medium). MW denotes the data in the Milky Way and Leo T is a gas-rich dwarf galaxy. We take the observed values of $T$ and $\dot{C}$ from Ref. [55].
@@ III Axions coupled to electromagnetic fields
As a warm-up, let us revisit the case where ${\bm{B}}={\bm{0}}$. Under the condition (23), we ignore the term $k^{2}{\bm{A}}$ relative to $\ddot{\bm{A}}$ in Eq. (22). On account of Eq. (24), we have ${\bm{J}}\simeq-\dot{\bm{E}}$ and $\dot{\bm{J}}\simeq\omega_{p}^{2}{\bm{E}}-\nu{\bm{J}}$ from Eqs. (22) and (25), respectively. Combining these two equations leads to
@@
$\omega_{c}=1.2\times 10^{-14}~{}{\rm eV}\left(\frac{B_{0}}{10^{-6}~{}{\rm G}}\right)\,,$ (36)
where we used the conversion of unit $1\,{\rm G}=1.95\times 10^{-2}$ eV2. For a typical magnetic field strength $B_{0}\approx 10^{-6}$ G, we have $\omega_{c}\simeq 10^{-14}$ eV. In such a case, so long as $n_{e}$ is not much less than the order $10^{-3}$ cm-3, the plasma frequency (7) is much larger than $\omega_{c}$, i.e., $\omega_{p}\gg\omega_{c}$. For $B_{0}\gtrsim 10^{-4}$ G, we have the opposite inequality $\omega_{p}\lesssim\omega_{c}$. Even for such large values of $B_{0}$, we are assuming the presence of a background magnetic field and an axion field which are both homogeneous, so that the local density of electrons is not significantly perturbed by them. Then, the approximation of an incompressible plasma for electrons ($\nabla\cdot{\bm{u}}_{e}=0$) does not lose its validity.
@@
The amplitudes of $A_{x}$ and $E_{x}$ blow up around the axion mass $m_{\phi}=\omega_{p}$, which is a characteristic feature of forced resonance. At $m_{\phi}=\omega_{p}$, the squared amplitude of $A_{x}$ has a peak value ${\cal D}_{\rm M}^{2}=(g\bar{\phi}_{0}B_{0})^{2}(\omega_{p}^{2}+\nu^{2})/(\omega_{p}^{2}\nu^{2})$. The region in which ${\cal D}^{2}\geq{\cal D}_{\rm M}^{2}/2$ corresponds to $\omega_{p}-\nu/2\leq m_{\phi}\leq\omega_{p}+\nu/2$, so that the resonance width is $\Delta m_{\phi}=\nu$. Since we are now considering the case $\nu\ll\omega_{p}$, ${\cal D}^{2}$ has a sharp peak around $m_{\phi}=\omega_{p}\pm\nu/2$. The length (or time) scale associated with the resonance width $\Delta m_{\phi}=\nu$ corresponds to $t_{d}=1/\nu$. Since $\nu\approx 10^{-22}$ eV for WNM (Leo T), we have $t_{d}\approx 10^{15}~{}{\rm m}\approx 0.1$ pc. On the other hand, the typical scale for the spatial variation of $n_{e}$ in the galactic plasma is of order $t_{e}\approx 0.1\sim 1$ kpc, see e.g., Ref. [65]. Since $t_{d}\ll t_{e}$, the resonance position $m_{\phi}=\omega_{p}=\sqrt{n_{e}e^{2}/m_{e}}$ hardly changes during the time interval $t_{d}$. Thus, so long as $t_{d}\ll t_{e}$, the efficiency of resonance is not spoiled by the spatial variation of $n_{e}$.
@@ IV Heating rate induced by forced resonance
Taking the homogeneous limit $k\to 0$ in Eq. (37) and using the property $E_{x}\simeq-\dot{A}_{x}$, we have
@@
We note that this has been derived by taking the large-scale limit $k\to 0$ in Fourier space, which amounts to neglecting all the spatial derivatives in real-space Eqs. (12)-(14). Since we are interested in the large-scale magnetic field whose wavelength is of order 100 pc or more, this translates to the wavenumber $k\lesssim 10^{-25}$ eV. We are interested in the ranges of $\omega_{p}$ and $m_{\phi}$ spanning around $10^{-12}$ eV, in which case $k\ll\{\omega_{p},m_{\phi}\}$. In Eqs. (39) and (40), the $k^{2}$ term is suppressed relative to $\omega_{p}^{2}$ and $m_{\phi}^{2}$ for the large-scale magnetic field of our interest. Hence it is a good approximation to take the limit $k\to 0$ for the estimation of $E_{x}$. Then, we can interpret $\rho_{x}=E_{x}^{2}/2+\omega_{p}^{2}A_{x}^{2}/2$ as the $x$-component of the photon energy density in real space obtained by keeping the time derivatives alone. The temporal change of $\rho_{x}$ averaged over the oscillating period of the axion $T_{\phi}=2\pi/m_{\phi}$ is given by
@@
$g\leq 1.9\times 10^{-14}~{}{\rm GeV}^{-1}\left(\frac{B_{0}}{10^{-6}~{}{\rm G}}\right)^{-1}\left(\frac{\nu}{10^{-22}~{}{\rm eV}}\right)^{1/2}\left(\frac{\rho_{\rm DM}}{0.3~{}{\rm GeV}\,{\rm cm}^{-3}}\right)^{-1/2}\left(\frac{\dot{C}}{10^{-27}~{}{\rm erg}~{}{\rm cm}^{-3}~{}{\rm s}^{-1}}\right)^{1/2}\,,$ (58)
which is valid at the resonance point with the axion mass $m_{\phi}=\omega_{p}$. For $B_{0}=10^{-6}$ G, $\nu=10^{-22}$ eV, $\rho_{\rm DM}=0.3$ GeV cm-3, and $\dot{C}=10^{-27}~{}{\rm erg}~{}{\rm cm}^{-3}~{}{\rm s}^{-1}$, we have the forecast constraint $g\leq 1.9\times 10^{-14}$ GeV-1. This is tighter than the observational bound $g\leq 5.4\times 10^{-12}$ GeV-1 extracted from polarization measurements of the MWD stars for the mass range $m_{\phi}\leq 3\times 10^{-7}$ eV [35]. Should we obtain tighter observational bounds on the cooling rate $\dot{C}$, the constraint on $g$ becomes more stringent. For larger $B_{0}$, the upper limit on $g$ also decreases in proportion to $B_{0}^{-1}$.
Figure 1: The coupling constant $g$ versus the axion mass $m_{\phi}$ for $\dot{Q}$ equivalent to the two cooling rates $\dot{C}=10^{-27}$ erg cm-3 s-1 (left) and $\dot{C}=10^{-31}$ erg cm-3 s-1 (right). In the left and right panels, we choose the background magnetic field strength to be $B_{0}=10^{-6}$ G and $B_{0}=1$ G, respectively, with $\omega_{p}=10^{-12}$ eV and $\rho_{\rm DM}=0.3\,{\rm GeV}$ cm-3 in both cases. The thick colored lines correspond to (a) $\nu=10^{-13}\omega_{p}$, (b) $\nu=10^{-10}\omega_{p}$, and (c) $\nu=10^{-7}\omega_{p}$, respectively. For decreasing $\nu$, the sharper resonance occurs with smaller minimum values of $g$. The dashed black lines represent the observational bound $g\leq 5.4\times 10^{-12}$ GeV-1 derived from polarization measurements of the MWD stars for the axion mass range $m_{\phi}\leq 3\times 10^{-7}$ eV [35].
@@
In Fig. 1, we plot $g_{\rm max}$ versus $m_{\phi}$ for three different values of $\nu$. We choose two combinations of $\dot{C}$ and $B_{0}$ in the left and right panels. In both panels, the plasma frequency and the dark matter density are fixed to be $\omega_{p}=10^{-12}$ eV and $\rho_{\rm DM}=0.3\,{\rm GeV}$ cm-3. For a given $\nu$, $g_{\rm max}$ has a minimum at the resonance mass $m_{\phi}=\omega_{p}$. For decreasing $\nu$, the minimum values of $g_{\rm max}$ get smaller. In both the two asymptotic mass regions $m_{\phi}\ll\omega_{p}$ and $m_{\phi}\gg\omega_{p}$, we can confirm that $g_{\rm max}$ tends to increase for smaller $\nu$.